$SPAD/src/input e04dgf.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1994.

———-

— * —

\getchunk{license}
)clear all
showArrayValues true
showScalarValues true
n:=2
es:=1.0
fu:=0.4373903597E-14
it:=50
lin:=0.9
list:=true
ma:=1.0E+20
op:=3.26E-12
pr:=10
sta:=1
sto:=2
ve:=0
x:Matrix SF:=
 [[-1.0 ,1.0 ]]
ef:Expression Float:=_
  (exp(X[1])*(4*X[1]**2+2*X[2]**2+4*X[1]*X[2]+2*X[2]+1))::EXPR FLOAT
objfun:ASP49(OBJFUN):= retract ef
result:=e04dgf(n,es,fu,it,lin,list,ma,op,pr,sta,sto,ve,x,-1,objfun)

———-

References^{[1] nothing}

4